OpenStudy (anonymous):

Need help with Algebra 2!!!

OpenStudy (anonymous):

Help me simplify this \[\frac{ 3-\frac{ 3 }{ x+1 } }{ \frac{ 3 }{ x+1 } }\]

OpenStudy (jhannybean):

first change \(3\) so that it has a common denominator with \(-\dfrac{3}{x+1}\)

OpenStudy (jhannybean):

that means multiplying both top and bottom of \(\dfrac{3}{1}\) by \(x+1\)

OpenStudy (anonymous):

Then wouldn't the 3 be 3x+3?

OpenStudy (jhannybean):

Over what?

OpenStudy (jhannybean):

\[\frac{3}{1} \cdot \frac{x+1}{x+1}\]

OpenStudy (misty1212):

\[\frac{ 3-\frac{ 3 }{ x+1 } }{ \frac{ 3 }{ x+1 } }\times \frac{x+1}{x+1}\]

OpenStudy (jhannybean):

That works too.

OpenStudy (misty1212):

yeah i like that way best cancel cancel cancel\[\cancel{}\]

OpenStudy (anonymous):

How did you know that you were supposed to multiply by x+1?

OpenStudy (jhannybean):

Because one number is subtracting a fraction, in order to combine them, you must turn the number into a fraction

OpenStudy (jhannybean):

Or simply cancel out the fraction, as @misty1212 showed us, \(\left(3 -\dfrac{3}{x+1}\right)\cdot x+1\)

OpenStudy (jhannybean):

Using this method, treat \((x+1)\) as like a number, and use the distributive property cancel out the fraction, and multiply the number.

OpenStudy (anonymous):

Alright, thanks!!!

OpenStudy (jhannybean):

No problem :D